# Homogeneous braids are visually prime

### v1

@@ 3.2. Murasugi sums and representing fixed essential arcs therein
Let $(\Sigma,\phi)$ be the essential Murasugi sum of $(\Sigma_{1},\phi_{1})$ and $(\Sigma_{2},\phi_{2})$ along a polygon $P$. Assume there is an essential arc $\gamma\subset\Sigma$ fixed by $\phi=\phi_{2}\circ\phi_{1}$. Then one may isotope $\phi_{1},\phi_{2}$$\gamma$, with an isotopy of $\phi_{i}$ supported in $\Sigma_{i}$, such that $\phi_{1}(\gamma)=\phi_{2}^{-1}(\gamma)$, and $\phi_{1}(\gamma)=\phi_{2}^{-1}(\gamma)=\gamma$ on $\Sigma\setminus P$.
@@
Figure 6. Left: the surface $\Sigma$ with the curves (green, labeled 1, 2, 3; and pink, labeled 4 and 5) of the Dehn twists factors of $\phi$, and the plumbing region $P$ (grey). Explicitly, $\Sigma$ arises from the depicted plane surface by identifying the oriented circles with an arrow head marking and by completing with a one-handle where indicated with square brackets. The curves and the plumbing regions are completed accordingly.
@@
The $3$–manifold associated with $(\Sigma,\phi)$ is diffeomorphic to $S^{3}$ since $(\Sigma,\phi)$ is equivalent to the open book constructed as follows. Start with an annulus with a positive Dehn twist (corresponding to a positive Hopf band in $S^{3}$), this twist corresponds to $\tau_{1}$. Then do the positive stabilizations corresponding to $\tau_{2}$ and $\tau_{3}$, and finally the negative stabilization corresponding to $\tau_{4}^{-1}$ to obtain $(\Sigma_{1},\phi_{1})$. It turns out that the binding of $(\Sigma_{1},\phi_{1})$ is the prime fibered knot $6_{2}$, while the binding of $(\Sigma,\phi)$, the result after the final plumbing with $(\Sigma_{2},\phi_{2})$, is the connected sum of the prime fibered knot $8_{21}$ and a negative Hopf link. Because of this, we think of this example as a sibling to Example 1.2 from the introduction, and invite the reader to explicitly realize the pages and the binding in $S^{3}$ to check that indeed the resulting bindings are $6_{2}$ and the connected sum of a Hopf link with $8_{21}$. Vice versa, we invite the reader to realize Example 1.2 as an open book with monodromy $\phi=\phi_{2}\circ\phi_{1}$ given by a product of Dehn twists and to realize $\phi_{1}$ and $\phi_{2}$ such that $\phi_{1}(\gamma)$ and $\phi_{2}^{-1}(\gamma)$ coincide and only differ from $\gamma$ in the plumbing region $P$.
@@
Overall, we have that $\delta_{1}$ and $\delta_{2}$ are isotopic, they both intersect all $s_{i}$ minimally, and their intersections with $s_{i}$ coincide for all $i$. By Lemma 3.6 below, there exists and isotopy of $\Sigma$ sending $\delta_{1}$ to $\delta_{2}$, fixing all $s_{i}$ pointwise. This isotopy can be written as composition of an isotopy with support in $\Sigma_{1}$ and an isotopy with support in $\Sigma_{2}\setminus P$. We now change $\phi_{1}$ and $\phi_{2}$ by composing with the first and second of those isotopies, respectively. In this way, we have achieved $\delta_{1}=\delta_{2}$. Moreover, observe the following. Since $\phi_{1}$ is the identity on $\Sigma\setminus\Sigma_{1}$, we have that $\delta_{1}=\gamma$ on $\Sigma\setminus\Sigma_{1}$. Similarly $\delta_{2}=\gamma$ on $\Sigma\setminus\Sigma_{2}$. This implies $\delta_{1}=\delta_{2}=\gamma$ on $\Sigma\setminus P$ as desired. ∎
@@ 3.3. Right-veering and left-veering for arcs and monodromies, and veering in polygons
Let $D$ be a disk, $A$ and $B$ finite sets of chords with $a\cap a^{\prime}=b\cap b^{\prime}=\varnothing$ for $a\neq a^{\prime}\in A$, $b\neq b^{\prime}\in B$, and such that $\bigcup_{a\in A}\partial a=\bigcup_{b\in B}\partial b$. Then either $A=B$, or there exists $a_{1},a_{2}\in A$, $b_{1},b_{2}\in B$ with $a_{i}\cap b_{i}=\{p_{i}\}\in\partial D$, and seen from $p_{1}$, $b_{1}$ is to the right of $a_{1}$, and seen from $p_{2}$, $b_{2}$ is to the left of $a_{2}$.
Moreover, if $D$ is an $n$–gon, and for each each chord $c\in A\cup B$, $c$ has endpoints on different sides, we can find chords as above with the additional condition that for each $i$, the other endpoint of $a_{i}$ is on a different side than the other endpoint of $b_{i}$.
Proof.
Assume towards a contradiction that $A\neq B$, and that for all $a\in A$, $b\in B$ such that $a\cap b=\{p\}\in\partial D$, $b$ is to the right of $a$ at $p$. Since $A\neq B$, there exist such $a,b$. Now, let $p^{\prime}$ be the other endpoint of $b$. Let $a^{\prime}$ be the chord in $A$ with $p^{\prime}\in a^{\prime}$. Since $p\notin a^{\prime}$, $b$ has to be to the right of $a^{\prime}$ as seen from $p^{\prime}$. Let $p^{\prime\prime}$ be the other endpoint of $a^{\prime}$. Let $b^{\prime}\in B$ be the chord such that $p^{\prime\prime}\in b^{\prime}$. See Figure 9.
@@
Now suppose that $D$ is an $n$–gon, and suppose no chord joins points on the same side of $\partial D$. Assume towards a contradiction that for any $a\in A,b\in B$ such that $a\cap b=\{p\}$, and $b$ is to the right of $a$, the other endpoints of $a$ and $b$ lie on the same side of $D$. Take one such pair $a,b$ (it exists by the previous paragraph). Let $p^{\prime}$ be the other endpoint of $b$, and $a^{\prime}$ the chord of $A$ with endpoint on $p^{\prime}$. Now, seen from $p^{\prime}$, $b$ is to the right of $a^{\prime}$ (if it were not, since $a\cap a^{\prime}=\varnothing$, $a^{\prime}$ would have both its endpoints on the same side of $D$, see the left hand side of Figure 10). However, this means that the other endpoint of $a^{\prime}$, call it $p^{\prime\prime}$, lies on the same side of $D$ as $p$. Then, let $b^{\prime}\in B$ be the chord with endpoint on $p^{\prime\prime}$. It must be to the right of $a^{\prime}$ seen from $p^{\prime\prime}$. Hence by the hypothesis, the other endpoint of $b^{\prime}$ must be on the same side of $D$ as $p^{\prime}$. We end up in a situation like in the previous case, see the right hand side of Figure 10.
@@
As a concrete example of this we can see on the bottom of Figure 6 the 4–gon $P$, where $A=\gamma\cap P$ and $B=\phi_{1}(\gamma)\cap P=\phi_{2}^{-1}(\gamma)\cap P$. At $p_{1}$ the chord of $B$ is to the right of the chord of $A$ (and ends on a different side), while at $p_{5}$ the chord of $B$ is to the left of the chord of $A$ (and also ends on a different side).
@@
If (after an isotopy of $\phi_{1}$ supported in $P$) $\phi_{1}(\gamma)=\gamma$ in all of $\Sigma$, then we also have $\phi_{2}^{-1}(\gamma)=\gamma$ in all of $\Sigma$ (after an isotopy of $\phi_{2}$ supported in $P$). Consider $\gamma\cap\Sigma_{2}$, which is a finite union of proper arcs in $\Sigma_{2}$, each of them fixed by $\phi_{2}^{-1}$. However, $\phi_{2}^{-1}$ is strictly left-veering (viewed as a map on $\Sigma_{2}$), hence each of those arcs is boundary parallel in $\Sigma_{2}$. Using that $\gamma$ and $P$ are essential and intersect minimally, one may deduce that $\gamma\cap\Sigma_{2}=\varnothing$. Thus, we have $\gamma\subset\Sigma_{1}\setminus P$, contradicting the assumption that $(\Sigma_{1},\phi_{1})$ has no fixed essential arcs disjoint from $P$.
Therefore, we may and do assume that the arc $\phi_{1}(\gamma)=\phi_{2}^{-1}(\gamma)$ does not agree with $\gamma$, even after any isotopy supported in $P$ (recall that they do agree outside of $P$). In other words, thinking of $P$ as a polygon in $\mathbb{R}^{2}$, the following two set of chords of $P$ differ: the set of chords $A$ given by straightening the components of $\gamma\cap P$ and the set of chords $B$ given by straightening the components of $\phi_{1}(\gamma)\cap P$. Observe that $A$ and $B$ satisfy the hypotheses of Lemma 3.7, including the $n$–gon condition because of minimal intersection of $\gamma$ and $\phi_{1}(\gamma)$ with the sides of $P$. Hence, Lemma 3.7 yields chords $a\in A$, $b\in B$ such that $a\cap b=\{p\}$, $b\geq a$ as seen from $p$, and the other endpoints of $a$ and $b$ (call them $p_{a}$ and $p_{b}$) lie on different sides of $P$. Let $i\in\{1,2\}$ such that $p\in\partial\Sigma_{i}$. Let $a^{\mathrm{ex}}$ be the unique subarc of $\gamma$ that starts at $p$ and is a proper arc in $\Sigma_{i}$. ($a^{\mathrm{ex}}$ extends the arc in $P$ corresponding to $a$, and may in fact just be equal to it, namely if $p_{a}\in\partial\Sigma_{i}$). Similarly extend $b$ to $b^{\mathrm{ex}}$. Then $b^{\mathrm{ex}}=\phi_{1}(a^{\mathrm{ex}})$ or $b^{\mathrm{ex}}=\phi_{2}^{-1}(a^{\mathrm{ex}})$ if $i=1$ or $i=2$, respectively. Since $b^{\mathrm{ex}}>a^{\mathrm{ex}}$ and $\phi_{2}^{-1}$ is left-veering, it follows that $i=2$ cannot happen. Therefore, we have $i=1$.
@@ 4. Iterative Murasugi summing of strictly veering open books
We define a construction of iterative Murasugi plumbings, encoded by a tree with vertices open books and edges indicating how they are plumbed together, that yields a class of prime open books, in fact open books without essential fixed arcs. The resulting open books subsume the links that arise from homogeneous braid diagrams without decomposition circles and all fibered arborescent links in $S^{3}$.
@@ 4.1. Figure-eight plumbings
Let $w=(\Sigma,\phi)$ be an open book and let $a$ be an arc in $\Sigma$. We describe a tree of open books $T$ with a growing such that the resulting open book is the figure-eight plumbing of $w$ along $a$. For this let $p$ and $n$ be open books corresponding to a positive and negative Hopf link, respectively, that is, their pages are an annulus, say $A_{p}$ and $A_{n}$, respectively, and their monodromies are a right-veering and left-veering Dehn twist, respectively. Let $T$ be the tree with vertices $w,p,n$ and edges $\{w,p\},\{p,n\}$. We pick the unique growing $\mathrm{gr}$ of $T$ that starts with $w$. We further pick summing regions $P$ and $P^{\prime}$ as decoration for $\{w,p\}$ and $\{p,n\}$, respectively as follows: both are $4$–gons, $P\subset\Sigma$ is a neighborhood of $a$, $P,P^{\prime}\subset A_{p}$ are disjoint and are neighborhoods of the cocore of $A_{p}$ and $P^{\prime}\subset A_{n}$ is a neighborhood of a cocore of $A_{p}$.
@@ 4.3. Primeness result
Let $T$ be a tree of open books, where each open book is strictly left-veering or strictly right-veering. Then the associated open book, well-defined up to conjugacy, has no fixed essential arcs; in particular, the fiber surface is a prime surface in the associated $3$–manifold.
@@
Next we choose a growing of $T$ such that whenever we have grown a vertex that lies in a $T_{i}$ or we start with a vertex that lies in $T_{i}$, for some $1\leq i\leq k$, we grow the rest of the vertices that lie in $T_{i}$ before other vertices are grown. We prove that the the open book associated to this tree with this growing has no essential fixed arcs.

### v2

@@ 3.2. Murasugi sums and representing fixed essential arcs therein
Let $(\Sigma,\phi)$ be the essential Murasugi sum of $(\Sigma_{1},\phi_{1})$ and $(\Sigma_{2},\phi_{2})$ along a polygon $P$. Assume there is an essential arc $\gamma\subset\Sigma$ fixed by $\phi=\phi_{2}\circ\phi_{1}$. Then one may isotope $\phi_{1},\phi_{2}$$\gamma$, with an isotopy of $\phi_{i}$ supported in $\Sigma_{i}$, such that $\phi_{1}(\gamma)=\phi_{2}^{-1}(\gamma)$ on $\Sigma$, and $\phi_{1}(\gamma)=\phi_{2}^{-1}(\gamma)=\gamma$ on $\Sigma\setminus P$.
@@
Figure 6. Left: the surface $\Sigma$ with the curves (green, labeled 1, 2, 3; and pink, labeled 4 and 5) of the Dehn twists factors of $\phi$, and the plumbing region $P$ (grey). Explicitly, $\Sigma$ arises from the depicted plane surface (red) by identifying the oriented circles with an arrow head marking and by completing with a one-handle where indicated with square brackets. The curves and the plumbing regions are completed accordingly.
@@
The 33–manifold associated with $(\Sigma,\phi)$ is diffeomorphic to $S^{3}$ since $(\Sigma,\phi)$ is equivalent to the open book constructed as follows. Start with an annulus with a positive Dehn twist (corresponding to a positive Hopf band in $S^{3}$), this twist corresponds to $\tau_{1}$. Then do the positive stabilizations corresponding to $\tau_{2}$ and $\tau_{3}$, and finally the negative stabilization corresponding to $\tau_{4}^{-1}$ to obtain $(\Sigma_{1},\phi_{1})$. It turns out that the binding of $(\Sigma_{1},\phi_{1})$ is the prime fibered knot $6_{2}$, while the binding of $(\Sigma,\phi)$, the result after the final plumbing with $(\Sigma_{2},\phi_{2})$, is the connected sum of the prime fibered knot $8_{21}$ and a negative Hopf link. Because of this, we think of this example as a sibling to Example 1.2 from the introduction, and invite the reader to explicitly realize the pages and the binding in $S^{3}$ to check that indeed the resulting bindings are $6_{2}$ and the connected sum of a Hopf link with $8_{21}$. Vice versa, we invite the reader to realize Example 1.2 as an open book with monodromy $\phi=\phi_{2}\circ\phi_{1}$ given by a product of Dehn twists and to realize $\phi_{1}$ and $\phi_{2}$ such that $\phi_{1}(\gamma)$ and $\phi_{2}^{-1}(\gamma)$ coincide and only differ from $\gamma$ in the plumbing region $P$.
@@
Overall, we have that $\delta_{1}$ and $\delta_{2}$ are isotopic, they both intersect all $s_{i}$ minimally, and their intersections with $s_{i}$ coincide for all $i$. By Lemma 3.6 below, there exists an isotopy of $\Sigma$ sending $\delta_{1}$ to $\delta_{2}$, fixing all $s_{i}$ pointwise. This isotopy can be written as composition of an isotopy with support in $\Sigma_{1}$ and an isotopy with support in $\Sigma_{2}\setminus P$. We now change $\phi_{1}$ and $\phi_{2}$ by composing with the first and second of those isotopies, respectively. In this way, we have achieved $\delta_{1}=\delta_{2}$. Moreover, observe the following. Since $\phi_{1}$ is the identity on $\Sigma\setminus\Sigma_{1}$, we have that $\delta_{1}=\gamma$ on $\Sigma\setminus\Sigma_{1}$. Similarly $\delta_{2}=\gamma$ on $\Sigma\setminus\Sigma_{2}$. This implies $\delta_{1}=\delta_{2}=\gamma$ on $\Sigma\setminus P$ as desired. ∎
@@ 3.3. Right-veering and left-veering for arcs and monodromies, and veering in polygons
Let $D$ be a disk, $A$ and $B$ finite sets of chords with $a\cap a^{\prime}=b\cap b^{\prime}=\varnothing$ for $a\neq a^{\prime}\in A$, $b\neq b^{\prime}\in B$, and such that $\bigcup_{a\in A}\partial a=\bigcup_{b\in B}\partial b$. Then either $A=B$, or there exist $a_{1},a_{2}\in A$, $b_{1},b_{2}\in B$ with $a_{i}\cap b_{i}=\{p_{i}\}\in\partial D$, and seen from $p_{1}$, $b_{1}$ is to the right of $a_{1}$, and seen from $p_{2}$, $b_{2}$ is to the left of $a_{2}$.
Moreover, if $D$ is an nn–gon, and for each chord $\in A\cup B, cc$ has endpoints on different sides, we can find chords as above with the additional condition that for $i=1,2$, the other endpoint of $a_{i}$ is on a different side than the other endpoint of $b_{i}$.
As an illustration of Lemma 3.7, consider for example the 4–gon $P$ on the bottom of Figure 6, where $A=\gamma\cap P$ and $B=\phi_{1}(\gamma)\cap P=\phi_{2}^{-1}(\gamma)\cap P$. At $p_{1}$ the chord of $B$ is to the right of the chord of $A$ (and ends on a different side), while at $p_{5}$ the chord of $B$ is to the left of the chord of $A$ (and also ends on a different side).
Proof of Lemma 3.7.
For the first claim, assume towards a contradiction that $A\neq B$, and that for all $a\in A$, $b\in B$ such that $a\cap b=\{p\}\in\partial D$, $b$ is to the right of $a$ at $p$. Since $A\neq B$, there exist such $a,b$. Now, let $p^{\prime}$ be the other endpoint of $b$. Let $a^{\prime}$ be the chord in $A$ with $p^{\prime}\in a^{\prime}$. Since $p\notin a^{\prime}$, $b$ has to be to the right of $a^{\prime}$ as seen from $p^{\prime}$. Let $p^{\prime\prime}$ be the other endpoint of $a^{\prime}$. Let $b^{\prime}\in B$ be the chord such that $p^{\prime\prime}\in b^{\prime}$. See Figure 9.
@@
For the second claim, suppose that $D$ is an nn–gon, and suppose no chord joins points on the same side of $\partial D$. Assume towards a contradiction that for any $a\in A,b\in B$ such that $a\cap b=\{p\}\in\partial D$, and $b$ is to the right of $a$, the other endpoints of $a$ and $b$ lie on the same side of $D$. Note that a pair $a,b$ where $b$ is to the right of $a$ at $a\cap b=\{p\}\in\partial D$ exists by the first claim, proved in the previous paragraph.
Then, take one such pair $a,b$. Let $p^{\prime}$ be the other endpoint of $b$, and $a^{\prime}$ the chord of $A$ with endpoint on $p^{\prime}$. Now, seen from $p^{\prime}$, $b$ is to the right of $a^{\prime}$ (if it were not, since $\cap a^{\prime}=\varnothing, aa^{\prime}$ would have both its endpoints on the same side of $D$, see the left hand side of Figure 10). However, this means that the other endpoint of $a^{\prime}$, call it $p^{\prime\prime}$, lies on the same side of $D$ as $p$. Then, let $b^{\prime}\in B$ be the chord with endpoint on $p^{\prime\prime}$. It must be to the right of $a^{\prime}$ seen from $p^{\prime\prime}$. Hence by the hypothesis, the other endpoint of $b^{\prime}$ must be on the same side of $D$ as $p^{\prime}$. We end up in a situation like in the previous case, see the right hand side of Figure 10.
@@
If (after an isotopy of $\phi_{1}$ supported in $P$) $\phi_{1}(\gamma)=\gamma$ in all of $\Sigma$, then we also have $\phi_{2}^{-1}(\gamma)=\gamma$ in all of $\Sigma$ (after an isotopy of $\phi_{2}$ supported in $P$). Consider $\gamma\cap\Sigma_{2}$, which is a finite union of proper arcs in $\Sigma_{2}$, each of them fixed by $\phi_{2}^{-1}$. However, $\phi_{2}^{-1}$ is strictly left-veering (viewed as a map on $\Sigma_{2}$), hence each of those arcs is boundary parallel in $\Sigma_{2}$. Using that $\gamma$ and $P$ are essential and $\gamma$ and $\partial P$ intersect minimally, one may deduce that $\gamma\cap\Sigma_{2}=\varnothing$. Indeed, $\gamma\cap\Sigma_{2}$ is a collection of boundary parallel arcs in $\Sigma_{2}$ (that is, together with segments of $\partial\Sigma_{2}$, they cut out bigons). Then, using these bigons we may isotope $\gamma$ freely along the boundary away from $\Sigma_{2}$. Thus, we have $\gamma\subset\Sigma_{1}\setminus P$, contradicting the assumption that $(\Sigma_{1},\phi_{1})$ has no fixed essential arcs disjoint from $P$ (since the Murasugi sum is essential, if $\gamma\subset\Sigma_{1}$ is essential in $\Sigma$ then it must also be essential in $\Sigma_{1}$).
Therefore, we may and do assume that the arc $\phi_{1}(\gamma)=\phi_{2}^{-1}(\gamma)$ does not agree with $\gamma$, even after any isotopy supported in $P$ (recall that they do agree outside of $P$). In other words, thinking of $P$ as a polygon in $\mathbb{R}^{2}$, the following two set of chords of $P$ differ: the set of chords $A$ given by straightening the components of $\gamma\cap P$ and the set of chords $B$ given by straightening the components of $\phi_{1}(\gamma)\cap P$. Observe that $A$ and $B$ satisfy the hypotheses of Lemma 3.7, including the nn–gon condition because of minimal intersection of $\gamma$ and $\phi_{1}(\gamma)$ with the sides of $P$. Hence, Lemma 3.7 yields chords $a\in A$, $b\in B$ such that $a\cap b=\{p\}$, $b\geq a$ as seen from $p$, and the other endpoints of $a$ and $b$ (call them $p_{a}$ and $p_{b}$) lie on different sides of $P$. Let $i\in\{1,2\}$ such that $p\in\partial\Sigma_{i}$. Let $a^{\mathrm{ex}}$ be the unique subarc of $\gamma$ that starts at $p$ and is a proper arc in $\Sigma_{i}$. ($a^{\mathrm{ex}}$ extends the arc in $P$ corresponding to $a$, and may in fact just be equal to it, namely if $p_{a}\in\partial\Sigma_{i}$). Similarly extend $b$ to $b^{\mathrm{ex}}$. Then $b^{\mathrm{ex}}=\phi_{1}(a^{\mathrm{ex}})$ or $b^{\mathrm{ex}}=\phi_{2}^{-1}(a^{\mathrm{ex}})$ if $i=1$ or $i=2$, respectively. Since $b^{\mathrm{ex}}>a^{\mathrm{ex}}$ and $\phi_{2}^{-1}$ is left-veering, it follows that $i=2$ cannot happen. Therefore, we have $i=1$.
@@ 4. Iterative Murasugi summing of strictly veering open books
We define a construction of iterative Murasugi sums, encoded by a tree with vertices open books and edges indicating how they are plumbed together, that yields a class of prime open books, in fact open books without essential fixed arcs. The resulting open books subsume the links that arise from homogeneous braid diagrams without decomposition circles and all fibered arborescent links in $S^{3}$.
@@ 4.1. Figure-eight plumbings
Let $w=(\Sigma,\phi)$ be an open book and let $a$ be an arc in $\Sigma$. We describe a tree of open books $T$ with a growing such that the resulting open book is the figure-eight plumbing of $w$ along $a$. For this let $p$ and $n$ be open books corresponding to a positive and negative Hopf link, respectively, that is, their pages are an annulus, say $A_{p}$ and $A_{n}$, respectively, and their monodromies are a right-veering and left-veering Dehn twist, respectively. Let $T$ be the tree with vertices $w,p,n$ and edges $\{w,p\},\{p,n\}$. We pick the unique growing $\mathrm{gr}$ of $T$ that starts with $w$. We further pick summing regions $P$ and $P^{\prime}$ as decoration for $\{w,p\}$ and $\{p,n\}$, respectively as follows: both are 44–gons, $P\subset\Sigma$ is a neighborhood of $a$, $P,P^{\prime}\subset A_{p}$ are disjoint and are neighborhoods of the cocore of $A_{p}$ and $P^{\prime}\subset A_{n}$ is a neighborhood of a cocore of $A_{p}$.
@@ 4.3. Primeness result
Let $T$ be a tree of open books, where each open book is strictly left-veering or strictly right-veering. Then the associated open book, well-defined up to conjugacy, has no fixed essential arcs; in particular, the fiber surface is a prime surface in the associated 33–manifold.
@@
Next we choose a growing of $T$ such that whenever we have grown a vertex that lies in a $T_{i}$ or we start with a vertex that lies in $T_{i}$, for some $1\leq i\leq k$, we grow the rest of the vertices that lie in $T_{i}$ before other vertices are grown. We prove that the open book associated to this tree with this growing has no essential fixed arcs.